Statistic problems an explanation would be appreciated

(5pts) Let X ,v N(25,3). Find the expected value, variance, and standard deviation of X. Find P(X 26.5) 5,3) Find the expected vale, variance, and standard d (5pts) Let XN (12,2). Find the expected value, variance, and standard deviation of X. Find P(x12.8). (Spts) Let Ybinomial(15,.65). Find the expected value, variance, and standard deviation of Y. Find P(Y 10)

(5pts) Let \(X \sim N(25,3)\). Find the expected value, variance, and standard deviation of

\(X .\) Find \(P(X \leq 26.5)\)

(5pts) Let \(X \sim N(12,2)\). Find the expected value, variance, and standard deviation of \(X\). Find \(P(X>12.8)\)

(5pts) Let \(Y \sim\) binomial \((15, .65)\). Find the expected value, variance, and standard deviation of \(Y\). Find \(P(Y=10)\).
Answer #1

Q1)

\(E(X)=\mu=25\)

\(V(X)=\sigma^{2}=9\)

\(S D(X)=\sigma=3\)

\(P(X \leq 26.5)=P\left(Z \leq \frac{26.5-25}{3}\right)\)

\(P(Z \leq 0.5)=P(-\infty

\(P(Z \leq 0.5)=0.5+0.1915=0.6915\)

Q 2)

\(E(X)=\mu=12\)

\(V(X)=\sigma^{2}=4\)

\(S D(X)=\sigma=2\)

\(P(X>12.8)=P\left(Z>\frac{12.8-12}{2}\right)\)

\(P(Z>0.4)=P(0

\(P(Z>0.4)=0.5-0.1554=0.3446\)

Q 3)

\(E(Y)=n p=15 * 0.65=9.75\)

\(V(Y)=n p q=15^{*} 0.65^{*} 0.35=3.4125\)

\(S D(Y)=\sqrt{3.4125}=1.8473\)

\(P(Y=10)=\left(\begin{array}{c}15 \\ 10\end{array}\right)(0.65)^{10}(0.35)^{15-10}=0.2123\)

Excel Command : =BINOM.DIST(10,15,0.65,FALSE)

> can u help me answer my quetion ?

1. Suppose that in a day, the probability of insurance agents not closing any deal is 0.3. On the other
hand, the probability the he can close deal is 0.25; two deal 0.35 and three deals, 0.1. Find the agents
expected number and variance of close deal in a day.
2. Find the expected number of correct answers and the variance in Quiz 2.

Czar Mohammad D. Musa Thu, Apr 15, 2021 1:48 AM